A point particle of charge 2.5 x 10^9 C and a mass of 3.65 x 10^-7 kg in uniform electric field directed to the right. It is released from rest and moves to the right. It is released from rest and moves to the right. After it has traveled 1.45 m, its speed is 35 m/s. Find,



a) work done on the particle



b) change in the electric potential energy of the particle

1
Expert's answer
2022-03-03T12:05:40-0500

"A=\frac{mv^2}2-mgh=0.22~mJ,"

"\Delta\varphi=\frac Aq=88\cdot 10^{-15}~V."